what would the next 2 numbers be for this sequence? 5, 15, 35, 65, 105
Answer by ewatrrr(24785)

5, 15, 35, 65, 105
Pattern:
a1 =5 = 5(2•0 + 1)
a2 = 15 = 5 (2•1 + 1)
a3 = 35 = 5(2•3 + 1)
a4 = 65 = 5(2*6 + 1_)
a5 = 105 = 5(2•10 + 1)
a6 = 5(2*15 + 1
a7 = 5(2*21 + 1)
etc